# Demystifying Spatial Confounding

### v2

@@ 5.4 Simulations for capped spatial+
where $\bm{\epsilon}^{y}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$ and $a$ measures the weight of the unconfounded high frequencies in $\mathbf{x}$. In order to generate $\mathbf{z}^{x}_{\mathrm{sp,low}}$, $\mathbf{z}^{x}_{\mathrm{sp,high}}$ and $\mathbf{z}^{x}_{\mathrm{sp,medium}}$, we fit a thin plate spline to a realization of a Gaussian process with $\kappa=0.5/3$ (see beginning of the section), where the penalty and design matrix have now been reparameterised such that they only consider the $100$ lowest, the $k$ highest and all medium frequencies, respectively; $\mathbf{z}^{z}_{\mathrm{sp,low}}$ is generated similarly to Section 4 but it is fitted to a new realization of a Gaussian process with $\kappa=0.5/3$ using only the $100$ lowest basis functions. We consider $k\in\{5,10,20\}$ and $a\in\{1,2\}$ in the data generating process; $\beta=1$, $\sigma=0.1$, $n=1000$, $N=50$. The reparametrisation is explained in Web Appendix 2.
@@
In the data analysis model, the capped spatial+ model uses a reparameterised $\tilde{\mathbf{B}}_{\mathrm{sp}}$ and $\tilde{\bm{\beta}}_{\mathrm{sp}}$ in the design and penalty matrix for the spatial effects in the first and second stage equation of spatial+ with $\mathrm{cap}\in\{5,10,15,20,25,30,40,50\}$, so that we consider cases where the used cap is above, below or equal to the true cap $k$. Unlike the standard spatial+, this allows us to control the frequencies included in the spatial effects for $\mathbf{y}$ and $\mathbf{x}$, such that ideally $\mathbf{r}^{x}$ should not have any confounded frequencies. Crucially here, on the first stage equation for $\mathbf{x}$ we use no smoothing in order to guarantee the residuals $\mathbf{r}^{x}$ are orthogonal to the subspace spanned by $\tilde{\mathbf{B}}_{\mathrm{sp}}\tilde{\bm{\beta}}_{\mathrm{sp}}$. We use $n$ basis functions in the spatial model.
@@
In Figure 3, the bias decreases for increasing $k$ and $a$, as larger $k$ implies less confounded medium frequencies and more unconfounded high frequencies, and larger $a$ implies more weight being put on unconfounded high frequencies in $\mathbf{x}$ and thus less bias. For $\mathrm{cap}\leq k$, we are able to completely remove confounding bias, even if only using a smaller fraction of the unconfounded high frequency component in $\mathbf{x}$ to identify the effect. When the $k<\mathrm{cap}$ the residuals in spatial+ contain medium frequency confounding and, quite noticeably, bias arises. Thus, we observe a constant bias of zero for caps below $k$ and a considerable increase for caps above $k$. The MSE for the predicted values reveals that in general capping increases MSE, although quite negligibly for $k\leq\mathrm{cap}$, when compared to the drastic increase observed for $k>\mathrm{cap}$ (see Web Appendix 3.1).
@@
Figure 3: Bias of $\widehat{\beta}$ in spatial model and capped spatial+ models with different combinations of $k\in\{5,10,20\}$ and $a\in\{1,2\}$. The true value is $\widehat{\beta}=1$. All other parameters remain constant.
@@ 6 Application to air temperature in Germany
As a motivating study, we analyse the relationship between the monthly mean temperature at two metres above ground, treated here as the response, and the monthly mean precipitation, the covariate, in Germany for the year 2010 at 336 locations. The data are open access and provided by the German Meteorological Service (DWD). For convenience, precipitation is measured in units of 10 millimetres and temperature is measured in degrees Celsius. In the context of spatial confounding, climate is a compelling example, as both covariates and response can exhibit relatively larger or narrower spatial correlation ranges, also depending on the time of year.
@@
All spatial effects are modelled with thin plate spline based on mgcv. The results are shown in Figure 4. As the cap used in our spatial+ increases, the fit of the model deteriorates, so we only show the estimates with $cap=5,6,\ldots,15$ (caps below 5 may not be reliable as there is likely little spatial information left in the residuals), which are also statistically significant at the 95% confidence level. From September to January and May, the caps (below 15) are generally insignificant, as indicated by the lack of pink dots in the plot. In the other months, the capped spatial+ can be used to assess whether we have confidence in the spatial estimate, and the dots show how the estimates should likely be adjusted. November is perhaps the most puzzling, where the coefficients appear to be either positive or negative depending on the frequencies included in the capped spatial+. On closer inspection, this can be explained by a more complex spatial structure of precipitation in this month, with many localised patterns, so that a linear effect of precipitation may not be sufficient in general. With the exception of November and December, all estimates plotted in Figure 4 are negative, indicating a generally negative association between the two variables.
@@
In order to understand the observed behaviour in more detail, we focus on three months: January, April and August. The results are shown in Table 1 for caps up to 9, which also shows the AIC and RMSE of the different models. We start with the spring and summer months. In April, the capped spatial+ estimates are lower in absolute value than those of the spatial model. This may indicate that the spatial model estimate is biased, and under the assumption that the highest frequencies are unconfounded, the estimate should be adjusted down (in absolute terms). In August, the estimates for all models appear to be fairly consistent, with little variability in the capped spatial+ estimates. This suggests that unobserved spatial confounders are not affecting the estimation and the spatial model estimate does not need to be adjusted. In January, we know from Figure  that none of the capped spatial+ estimates up to a cap of 15 are statistically significant. This suggests that the highest frequencies (up to 15) in the spatial effect are not helping us to identify the effect of rainfall and may simply not be present in this particular covariate. As mentioned in Section 5.2, however, when a covariate is dominated by low frequencies, the estimate in the spatial model may well be biased. In this case, we can adapt the capped spatial+ method to identify the highest frequencies at which $\mathbf{x}$ has a significant effect. Assuming that these frequencies are unconfounded, we can use the corresponding capped spatial+ estimate to assess the bias. We use a sliding window technique that systematically removes 15 consecutive frequencies, starting progressively deeper into the low frequency spectrum, where $start$ is the starting frequency. Figure 4 shows the estimates for the first two sliding windows, with significant estimates at the 5% level represented by triangles. We can see that in January the spatial model may overestimate the effect (in absolute terms). We followed the same approach for all the remaining autumn and winter months and for May. The results suggest that the extra peak building up to October may be an artefact, with values likely closer to flat around $-0.125$. Apart from the peak in April – perhaps a case of “Der April macht, was er will” (“April does as it pleases”) – the coefficients remain relatively stable. In all the months analysed using the sliding window technique, the first significant coefficient appears the latest at $start=8$, except in September, where it appears at $start=38$.
@@ 7 Discussion
As illustrated in this paper, the analytical expressions derived in Proposition 1 and Corollary 1 provide an intuitive framework for explaining and, indeed, demystifying the bias arising from spatial confounding. In particular, we show that bias in the spatial model is caused by spatial smoothing and can become arbitrarily large. In practical applications, the difference between the estimates in the spatial and non-spatial models is commonly seen as an indicator of spatial confounding. However, our analysis and simulations show that this difference is not directly related to the size or occurrence of bias in the spatial model. Instead we propose using the methods set out in Section 5 to assess and adjust for the bias.
As most other work on spatial confounding, the focus in this paper has been on quantifying the bias in spatial regressions. A derivation and short discussion about the variance of the covariate effect estimate is included in Web Appendix 1.6. However, to put the bias into perspective and thoroughly conduct statistical inference, more investigations on associated uncertainty will be an important direction for future research.
@@
The data is publicly available on the website of the German weather service (DWD).
Supporting Information
Web Appendices referenced in Section 26, are included below.
See pages - of appendix.pdf

### v3

@@ 5.4 Simulations for capped spatial+
where $\bm{\epsilon}^{y}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$ and $a>0$ controls the size of the unconfounded high frequency spatial component in $\mathbf{x}$. In order to generate $\mathbf{z}^{x}_{\mathrm{sp,low}}$, $\mathbf{z}^{x}_{\mathrm{sp,high}}$ and $\mathbf{z}^{x}_{\mathrm{sp,medium}}$, we fit a thin plate spline to a realization of a Gaussian process with $\kappa=0.5/3$ (see beginning of the section), where the penalty and design matrix have now been reparameterised such that they only consider the $100$ lowest, the $k$ highest and all medium frequencies, respectively; $\mathbf{z}^{z}_{\mathrm{sp,low}}$ is generated similarly to Section 4 but it is fitted to a new realization of a Gaussian process with $\kappa=0.5/3$ using only the $100$ lowest basis functions. We consider $k\in\{5,10,20\}$ and $a\in\{1,2\}$ in the data generating process; $\beta=1$, $\sigma=0.1$, $n=1000$, $N=50$. The reparametrisation is explained in Appendix 2.
@@
In the data analysis model, the capped spatial+ model uses a reparameterised $\tilde{\mathbf{B}}_{\mathrm{sp}}$ and $\tilde{\bm{\beta}}_{\mathrm{sp}}$ in the design and penalty matrix for the spatial effects in the first and second stage equation of spatial+ with $\mathrm{cap}\in\{5,10,15,20,25,30,40,50\}$, so that we consider cases where the used cap is above, below or equal to the true cap $k$. Unlike the standard spatial+, this allows us to control the frequencies included in the spatial effects for $\mathbf{y}$ and $\mathbf{x}$, such that ideally $\mathbf{r}^{x}$ should not have any confounded frequencies. Here, on the first stage regression for $\mathbf{x}$ we use no smoothing in order to guarantee the residuals $\mathbf{r}^{x}$ are orthogonal to the subspace spanned by $\tilde{\mathbf{B}}_{\mathrm{sp}}\tilde{\bm{\beta}}_{\mathrm{sp}}$. We use $n$ basis functions in the spatial model.
@@
In Figure 3 we see that the bias decreases for increasing $k$ and $a$. This is expected from our results as a larger $k$ implies fewer confounded medium frequencies and more unconfounded high frequencies, while a larger $a$ implies a larger proportion of unconfounded high frequencies in $\mathbf{x}$. For $\mathrm{cap}\leq k$, we are able to completely remove confounding bias, even if only using a smaller fraction of the unconfounded high frequency component in $\mathbf{x}$ to identify the effect. When the $k<\mathrm{cap}$ the residuals in spatial+ contain medium frequency confounding and, quite noticeably, bias arises. Thus, we observe a constant bias of zero for caps below $k$ and a considerable increase for caps above $k$. The MSE for the predicted values reveals that in general capping increases MSE, although quite negligibly for $k\leq\mathrm{cap}$, when compared to the drastic increase observed for $k>\mathrm{cap}$ (see Appendix 3.1).
@@
Figure 3: Bias of $\widehat{\beta}$ in the spatial and capped spatial+ models with different combinations of $k\in\{5,10,20\}$ and $a\in\{1,2\}$. The true value is $\beta=1$. All other parameters remain constant.
@@ 6 Application to air temperature in Germany
As an empirical illustration, we analyse the relationship between the monthly mean temperature at two metres above ground, treated here as the response, and the monthly mean precipitation, the covariate, in Germany for the year 2010 at 336 locations. The data are open access and provided by the German Meteorological Service (DWD). For convenience, precipitation is measured in units of 10 millimetres and temperature is measured in degrees Celsius. In the context of spatial confounding, climate is a compelling example, as both covariates and response can exhibit relatively larger or narrower spatial correlation ranges, also depending on the time of year.
@@
All spatial effects are modelled with thin plate regression splines in mgcv (Wood,, 2024). The results are shown in Figure 4. When the cap used for capped spatial+ is very small, estimates of the covariate effect may not be reliable, as the residuals will be very small and unlikely to have sufficient information about the covariate. On the other hand, as the cap increases, the fit of the model deteriorates. Therefore, we only show the estimates with $cap=5,6,\ldots,15$ which are also statistically significant at the 95% confidence level. From September to January and May, the caps (below 15) are generally insignificant, as indicated by the lack of pink dots in the plot. In the other months, the capped spatial+ estimates can be used to assess whether we have confidence in the spatial estimate, and the dots show how the estimate should likely be adjusted. November is perhaps the most puzzling, where the coefficients appear to be either positive or negative depending on the frequencies included in capped spatial+. On closer inspection, this can be explained by a more complex spatial structure of precipitation in this month, with many localised patterns, so that a linear effect of precipitation may not be sufficient in general. With the exception of November and December, all estimates plotted in Figure 4 are negative, indicating a generally negative association between the two variables.
@@
In order to understand the observed behaviour in more detail, we focus on three months: January, April and August. For these three months, results for caps up to 9, including the AIC and RMSE of the different models, are shown in Table 1. We start with the spring and summer months. In April, the capped spatial+ estimates are lower in absolute value than those of the spatial model. This may indicate that the spatial model estimate is biased, and under the assumption that the highest frequencies are unconfounded, the estimate should be adjusted down (in absolute terms). In August, the estimates for all models appear to be fairly consistent, with little variability in the capped spatial+ estimates. This suggests that unobserved spatial confounders are not affecting the estimation and the spatial model estimate does not need to be adjusted. In January, we know from Figure 4 that none of the capped spatial+ estimates up to a cap of 15 are statistically significant. This suggests that the highest frequencies (up to 15) in the spatial effect are not helping us to identify the effect of rainfall and may simply not be present in this particular covariate. As mentioned in Section 5.2, however, when a covariate is dominated by low frequencies, the estimate in the spatial model may well be biased. In this case, we can adapt the capped spatial+ method to identify the highest frequencies at which $\mathbf{x}$ has a significant effect. Assuming that these frequencies are unconfounded, we can use the corresponding capped spatial+ estimate to assess the bias. We use a sliding window technique that systematically removes 15 consecutive frequencies, starting progressively deeper into the low frequency spectrum, where $start$ is the starting frequency. Figure 4 shows the estimates for the first two sliding windows, with significant estimates at the 5% level represented by triangles. We can see that in January the spatial model may overestimate the effect (in absolute terms). We followed the same approach for all the remaining autumn and winter months and for May. The results suggest that the extra peak building up to October may be an artefact, with values likely closer to flat around $-0.125$. Apart from the peak in April – perhaps a case of “Der April macht, was er will” (“April does as it pleases”) – the coefficients remain relatively stable. In all the months analysed using the sliding window technique, the first significant coefficient appears the latest at $start=8$, except in September, where it appears at $start=38$.
@@ 7 Discussion
As illustrated in this paper, the analytical expressions derived in Proposition 1 and Corollary 1 provide an intuitive framework for explaining and, indeed, demystifying the bias arising from spatial confounding. In particular, we show that bias in the spatial model is caused by spatial smoothing and can become arbitrarily large. In practical applications, the difference between the estimates in the spatial and non-spatial models is commonly seen as an indicator of spatial confounding. However, our analysis and simulations show that a difference between these estimates does not in itself show whether the estimate in the spatial model is biased. Instead we propose using the methods set out in Section 5 to assess and adjust for the bias.
As most other work on spatial confounding, the focus in this paper has been on quantifying the bias in spatial regressions. A derivation and short discussion about the variance of the covariate effect estimate is included in Appendix 1.6. However, to put the bias into perspective and thoroughly conduct statistical inference, more investigations on associated uncertainty will be an important direction for future research.
@@
The data is publicly available on the website of the German weather service (DWD). We provide code for all analyses in the paper, which includes instructions on how to download the data.